OpenStudy (anonymous):

How would you show 0 vector is in K if it is K= (X is in M, bx=xb)?

OpenStudy (kinggeorge):

So \(K=\{x\in M|bx=xb\}\), where \(M\) is a vector space?

OpenStudy (kinggeorge):

And \(b\) is just a fixed vector in \(M\)?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

M is a subspace of a vector space actually..

OpenStudy (kinggeorge):

Well, what is \(b*0\) and what is \(0*b\)?

OpenStudy (anonymous):

can I just say that b is a 0 vector?

OpenStudy (kinggeorge):

No. That's not a good way to do it. But what is \(b*0\)?

OpenStudy (anonymous):

so just say that 0=0?

OpenStudy (anonymous):

how is vector product defined?

OpenStudy (kinggeorge):

That's something else I just recently noticed as well. Are these matrices?

OpenStudy (anonymous):

yes!

OpenStudy (kinggeorge):

So if you take any matrix and multiply it by the zero matrix, what do you get?

OpenStudy (anonymous):

0

OpenStudy (kinggeorge):

Right. So \(b*0=0\) and \(0*b=0\). So \(b*0=0*b\), and \(0\in K\).

OpenStudy (anonymous):

okay I get it now, thanks

OpenStudy (kinggeorge):

You're welcome.